### v1

# Block-to-Scene Pre-training for Point Cloud Hybrid-Domain Masked Autoencoders

Abstract
Point clouds, as a primary representation of 3D data, can be categorized into scene domain point clouds and object domain point clouds based on the modeled content. Masked autoencoders (MAE) have become the mainstream paradigm in point clouds self-supervised learning. However, existing MAE-based methods are domain-specific, limiting the model’s generalization. In this paper, we propose to pre-train a general Point cloud Hybrid-Domain Masked AutoEncoder (PointHDMAE) via a block-to-scene pre-training strategy. We first propose a hybrid-domain masked autoencoder consisting of an encoder and decoder belonging to the scene domain and object domain, respectively. The object domain encoder specializes in handling object point clouds and multiple shared object encoders assist the scene domain encoder in analyzing the scene point clouds. Furthermore, we propose a block-to-scene strategy to pre-train our hybrid-domain model. Specifically, we first randomly select point blocks within a scene and apply a set of transformations to convert each point block coordinates from the scene space to the object space. Then, we employ an object-level mask and reconstruction pipeline to recover the masked points of each block, enabling the object encoder to learn a universal object representation. Finally, we introduce a scene-level block position regression pipeline, which utilizes the blocks’ features in the object space to regress these blocks’ initial positions within the scene space, facilitating the learning of scene representations. Extensive experiments across different datasets and tasks demonstrate the generalization and superiority of our hybrid-domain model. The code will be released.
1 Introduction
With the rapid development of 3D scanning technology, 3D point clouds have become the mainstream representation for 3D objects due to their ease of acquisition, explicit representation, and efficient storage. Point clouds can be categorized into scene domain point clouds (Dai et al., 2017; Song et al., 2015; Armeni et al., 2016; Zheng et al., 2020; Sun et al., 2020) and object domain point clouds (Wu et al., 2015; Chang et al., 2015; Uy et al., 2019; Deitke et al., 2024; Yu et al., 2023) based on the modeling object. As shown in Figure 1 (a), object domain point clouds describe specific objects or entities, such as an airplane, with relatively fewer points. Scene domain point clouds represent the entire environment or scene, such as indoor scenes, including multiple objects, structures, and background elements, with a larger number of points. Due to the significant disparity in point count and the elements being described, a substantial domain gap exists in these two types of point clouds.
Figure 1: The handling approach for point clouds from different domains in domain-specific models (a) compared to our hybrid-domain model (b), and the generalization experiments of the domain-specific Point-MAE model (c).
Recently, point cloud masked autoencoders (Yu et al., 2022; Pang et al., 2022; Zhang et al., 2022; Dong et al., 2023; Zha et al., 2024a), pre-trained on massive point cloud data, have become the mainstream paradigm in point cloud self-supervised learning and have been widely applied to various point cloud tasks. It is inspired by masked image modeling (Bao et al., 2021; He et al., 2022; Xie et al., 2022), using the unmasked portions to predict the geometric coordinates or semantic features of the masked parts, thereby enabling the model to learn universal 3D representations. Despite the significant success, most of these methods are domain-specific due to the notable domain gap between object-domain and scene-domain point clouds. As shown in Figure 1 (a), these methods (Yu et al., 2022; He et al., 2022; Zhang et al., 2022; Zha et al., 2023b; Zhou et al., 2024) use scene-level models for scene tasks and object-level models for object tasks, thereby limiting their generalizability. For example, Point-MAE (Pang et al., 2022), which is pre-trained on ShapeNet (Chang et al., 2015), primarily performs object point cloud tasks. For scene point clouds, it requires re-pretraining on scene-level datasets like ScanNet (Dai et al., 2017) to adapt to the scene domain. As shown in Figure 1 (c), directly transferring an object domain pre-trained model (e.g.Point-MAE (Pang et al., 2022)) to scene tasks results in a significant performance drop. Similarly, models pre-trained on the scene domain also exhibit performance declines when transferred to object task.
Pre-training a general point cloud model is our persistent pursuit; however, it is highly challenging for two main reasons. Firstly, the input data is inconsistent. Scene-level point clouds, such as ScanNet (Dai et al., 2017), typically consist of 50k points, while object-level point clouds like ModelNet (Wu et al., 2015) typically consist of 1k points. The disparity in point count between the two types is significant, making it difficult to process both types of data simultaneously using a single model. Secondly, there is inconsistency in task emphasis. Scene point clouds typically involve object detection or segmentation, which often prioritizes understanding fine-grained local point clouds. On the contrary, object point clouds generally involve classification tasks, which tend to prioritize understanding global geometry.
To address the aforementioned challenges, we propose a block-to-scene pretraining strategy to pre-train a Point cloud Hybrid-Domain Masked Auto-Encoder (PointHDMAE). We address the challenge of inconsistent input data by using domain-specific encoders to process data from their respective domains. Additionally, we finetune the pre-trained model to address the inconsistency of task emphasis. Specifically, as shown in Figure 2, we first design a point cloud hybrid-domain architecture that consists of an encoder and decoder belonging to the scene domain and object domain, respectively. In the fine-tuning phase, as shown in Figure 1 (b), for object domain data, our model selectively activates the object-domain encoder for analysis. However, in the case of scene point clouds, we activate multiple shared object encoders to assist the scene encoder in analyzing scene domain data collaboratively.
Furthermore, we propose a block-to-scene pre-training strategy that couples masked reconstruction and position regression tasks of random object blocks within a scene for self-supervised learning, enabling us to train encoders for different domains simultaneously. Specifically, we first randomly select point blocks within a scene and apply a set of transformations to convert each point block coordinates from the scene space to the object space. Then, within the object domain, we use a mask and reconstruction pipeline to recover the masked points of each block, enabling it to learn universal object representations. Finally, we introduce a scene-level block position regression pipeline, which utilizes the blocks’ features in the object space to regress these blocks’ initial positions within the scene space, enabling the scene encoder to learn scene representations with the assistance of the object encoders. By block-to-scene pretraining, our model can simultaneously learn powerful object-level and scene-level representations and exhibit superior transferability. Our model can be fine-tuned directly on downstream tasks such as object point cloud classification, segmentation, completion, and scene point cloud detection without the need for any additional domain adaptation training.
Our main contributions can be summarized as follows: (1) We propose a point cloud hybrid-domain masked autoencoders to address the generalization limitations of existing domain-specific MAE-based models. (2) We propose a block-to-scene pretraining strategy, a joint pre-training strategy that reconstructs and regresses random object blocks within a scene. (3) Extensive experiments across different datasets and tasks demonstrate the generalization and superiority of our model.
2 Related Work
2.1 Self-supervised Learning for Point Cloud.
Self-supervised learning, which enables the learning of general representations from large amounts of unlabeled data, has been widely applied in fields such as language (Semnani et al., 2019; Brown et al., 2020; Achiam et al., 2023) and image (Bao et al., 2021; Chen et al., 2020b; a; He et al., 2022). Inspired by the success of visual pretraining, numerous point cloud pretraining methods have also been proposed. Based on the pretext tasks, they can be categorized into contrastive learning paradigms (Oord et al., 2018; Tian et al., 2020) and masked reconstruction paradigms (Bao et al., 2021; He et al., 2022). PointContrast (Xie et al., 2020b), CrossPoint (Afham et al., 2022), and DepthContrast (Zhang et al., 2021) construct positive and negative sample pairs using various methods and employ contrastive learning techniques to learn 3D representations.
Point-BERT (Yu et al., 2022) was the first to propose learning universal 3D representations using the paradigm of masked reconstruction. Subsequently, numerous explorations (Pang et al., 2022; Zhang et al., 2022; Zha et al., 2024a; b) have improved masked reconstruction from various perspectives. Point-MAE (Pang et al., 2022) and PointFEMAE (Zha et al., 2024a) introduced the masked autoencoder for reconstruction, and PointGPT (Chen et al., 2024) proposed pretraining using an autoregressive approach. To address the limited amount of point cloud during pretraining, many approaches integrate multimodal knowledge to aid in learning point cloud features. ACT (Dong et al., 2023) leverages a pre-trained image model to assist in point cloud reconstruction, while I2P-MAE (Zhang et al., 2023) employs an image-guided masking strategy. PiMAE (Chen et al., 2023) proposes to address the challenges of multi-modal interaction between point cloud and RGB image data through mask alignment, a two-branch MAE pipeline, and a cross-modal reconstruction module. In this paper, we propose a block-to-scene pretraining strategy that combines masked reconstruction and position regression in a joint self-supervised learning method to enhance the model’s generalizability.
3 Methodology
In this section, we provide a detailed explanation of how to use our block-to-scene pretraining strategy to train our point cloud hybrid-domain masked autoencoder.
3.1 Point Cloud Hybrid-Domain Masked Autoencoder (PointHDMAE)
The overall architecture of our PointHDMAE is shown in Figure 2, it is composed of four main components: a scene encoder, a scene decoder, a shared object encoder, and a shared object decoder. It is primarily used for two main task pipelines: object point cloud processing and scene point cloud processing.
Our PointHDMAE is a hybrid model. Due to significant differences across domains of point clouds and tasks, we selectively activate different sub-modules for various downstream data and tasks. For instance, in tasks such as object point cloud classification and object part segmentation, we selectively activate our object encoder according to specific tasks. For scene point cloud detection tasks, we activate all encoders. This collaborative approach is primarily adopted because utilizing the object encoder for analyzing local point blocks within the scene contributes to the scene encoder’s comprehension of scene intricacies.
3.2 Block-to-Scene Pretraining
Our block-to-scene pretraining primarily consists of the following three key components: random point block generation, object-level block masked reconstruction, and scene-level block position regression. Below, we provide a detailed illustration of the specific implementation of each component.
Figure 2: The architecture of our point cloud hybrid-domain masked autoencoder and the pipeline for block-to-scene pre-training. The left side illustrates the scene-level block position regression, while the right side shows the object-level block masked reconstruction.
3.2.1 Random Point Block Generation
Random point block selection. To leverage scene local details for scene understanding in an unsupervised manner, we randomly select $K_{o}$ local point blocks from the entire scene. We first random select $K_{o}$ points from the entire scene point cloud as the center points for each point block. For each center point, we then use the K-nearest neighbors algorithm to select the nearest $N_{o}$ points around it, forming the initial point block objects $\bm{B=\{B^{1},...,B^{K_{o}}\}}$, where the $i$-th point block is $\bm{B^{i}}\in\mathbb{R}^{N_{o}\times 3}$.
Ground-truth block position generation. We generate the ground truth position of each random point block in the scene, which will be used to constrain the predicted position regressed by the final scene decoder. Inspired by the detection (Carion et al., 2020; Misra et al., 2021; Dai et al., 2021) task, we use the 3D bounding box of each random point block as its ground truth position. By computing the mean of all points in each dimension of the entire point block, the coordinates of the center point are obtained. The half-lengths of the bounding box in each dimension are calculated by subtracting the minimum value from the maximum value in each dimension and dividing by 2. Subsequently, the center point coordinates and half-lengths in each dimension (x, y, and z) are concatenated to form the bounding box. Finally, standard procedures (Misra et al., 2021) are applied to compute bounding box parameters $\bm{B_{b}}$ such as size and corners for each bounding box.
3.2.2 Object-level Block Masked Reconstruction
Object point block generation. We treat each randomly selected point block in $\bm{B}$ as an object point block and transform its coordinates from the scene space to the object space for processing by the object autoencoder. Specifically, we apply a simple set of transformation functions to each point block. First, we subtract the coordinates of the center point from the $N_{o}$ local points to obtain the relative coordinates of each point. Then, we normalize these coordinates to the range [-1, 1]. Finally, after applying a random rotation transformation to each point block, we obtain all point blocks $\bm{B_{o}=\{B_{o}^{1},...,B_{o}^{K_{o}}\}}$ as input to the object encoder. Through these transformation functions, we convert the coordinates of each point block from the scene space to the object space, decoupling the point block object coordinates from the original scene coordinates. This enables the object encoder to learn the universal shape features of the point block objects.
Object point block masked and reconstruction. We use a shared object autoencoder to perform mask-based reconstruction self-supervised learning on all generated object point blocks $\bm{B_{o}}$, enabling our object encoder to learn a general representation of objects. We illustrate the entire mask and reconstruction process for the object point blocks using Point-MAE (Pang et al., 2022) as an example. For the $i$-th object point block $\bm{B_{o}^{i}\in\mathbb{R}^{N_{o}\times 3}}$, we use farthest point sampling and the K-nearest neighbors algorithm to divide it into $M_{o}$ point patches. Then, after randomly masking most of the patches, we generate initial tokens and positional encodings for each unmasked patch using MLP-based token encoding and positional encoding. By adding them, we obtain the token $\bm{E}_{o}^{0}\in\mathbb{R}^{rM_{o}\times C_{o}}$ for each unmasked patch, where $r$ represents the unmasked ratio, and $C_{o}$ denotes the object feature dimension. Finally, we use a shared object encoder to extract object features $\bm{E}_{o}^{n_{o}}\in\mathbb{R}^{rM_{o}\times C_{o}}$, where $n_{o}$ is the number of layers in the scene encoder.
In the decoding phase, we concatenate $\bm{E}_{o}^{n_{o}}$ with randomly initialized masked tokens to obtain $\bm{D}_{o}^{0}\in\mathbb{R}^{M_{o}\times C_{o}}$. Then, we use a standard Transformer-based decoder to decode, getting $\bm{D}_{o}^{m_{o}}\in\mathbb{R}^{M_{o}\times C_{o}}$. Finally, we use an MLP-based reconstruction head to reconstruct the coordinates of the masked point patches $\bm{R}_{o}^{i}\in\mathbb{R}^{N_{o}\times 3}$.
3.2.3 Scene-level Block Position Regression
Scene encoding. Given an input point cloud $\bm{P_{s}}\in\mathbb{R}^{N_{s}\times 3}$ with $N_{s}$ points, we first use farthest point sampling and the K-nearest neighbors algorithm to partition it into blocks. Then, using an MLP-based token encoding layer and a positional encoding layer, we generate the semantic token and positional encoding for each patch. By adding them, we obtain the token $\bm{E}_{s}^{0}\in\mathbb{R}^{M_{s}\times C_{s}}$ for each patch, where $M_{s}$ represents the number of scene patches, and $C_{s}$ denotes the scene feature dimension. Finally, we use a scene encoder based on the standard Transformer (Vaswani et al., 2017) architecture to extract scene features $\bm{E}_{s}^{n_{s}}\in\mathbb{R}^{M_{s}\times C_{s}}$, where $n_{s}$ is the number of Transformer layers in the scene encoder.
Scene decoding and position regression. We apply max pooling to the features of all blocks output by the object encoder to obtain the global feature for each block. After passing through the projection layer, these point block features are transformed into features $\bm{B}_{g}\in\mathbb{R}^{K_{o}\times C_{s}}$ for scene decoding input. However, we prevent the gradients of $\bm{B}_{g}$ from propagating backward into the mask reconstruction pipeline during the backpropagation process, thereby mitigating the multi-task interference caused by the scene regression task on the object encoder. We will provide a detailed explanation of this issue in Section 4.3.1. We then add the transformed point block features to randomly initialized queries to obtain enhanced queries $\bm{Q^{0}}\in\mathbb{R}^{q\times C_{s}}$, where $q$ is the number of queries. Since the number of point blocks and queries often differ, we replicate $\bm{B}_{g}$ to match all queries.
We use a Transformer decoder based on self-attention and cross-attention as our scene decoder. The input queries $\bm{Q^{0}}$, with the assistance of the encoded features $\bm{E}_{s}^{n_{s}}$, pass through our decoder to obtain the decoded query features $\bm{Q^{m_{s}}}$, where $m_{s}$ is the number of scene decoder. Finally, we use different MLP-based reconstruction heads to predict the 3D bounding box $\bm{B_{b}^{p}}$ of each random point block.
3.2.4 Loss Function.
We use a combination of mask reconstruction loss and point block regression loss to jointly constrain our pre-training process. For the reconstruction loss calculation, we follow previous work (Pang et al., 2022; Zhang et al., 2022) and use Chamfer Distance (Fan et al., 2017) (CD) as the loss function. For the regression loss calculation, we reference detection (Misra et al., 2021) and use Generalized Intersection over Union (Rezatofighi et al., 2019) (GIoU) as the loss function. Therefore, our loss function is defined as follows:
$\mathcal{L}=\lambda_{1}\cdot{CD}(\bm{R_{o}},\bm{B_{o}})+\lambda_{2}\cdot{GIoU}(\bm{B_{b}^{p}},\bm{B_{b}})$ (1)
where $\lambda_{1}$ and $\lambda_{2}$ is a weighted combination of reconstruction loss and regression loss.
4 Experiments
First, we pre-train the PointHDMAE model using our block-to-scene pretraining strategy based on point cloud data from the scene domain. After pre-training, we directly transfer the pre-trained model to various downstream tasks in different point cloud domains for fine-tuning. During fine-tuning, we selectively activate different sub-modules depending on the domain of the point cloud; for instance, we activate the object encoder for object point clouds, while utilizing all encoders for scene point clouds. This strategy enables our PointHDMAE model, pre-trained with the block-to-scene approach, to outperform existing domain-specific models in most cases without any additional domain adaptation training, demonstrating the generalization capability of our model.
4.1 Pre-training
Datasets. We combined all training data from the two most commonly used indoor scene datasets, SUNRGB-D (Song et al., 2015) and ScanNetV2 (Dai et al., 2017), to construct our pretraining dataset. Specifically, SUNRGB-D includes 5K single-view RGB-D training samples with oriented bounding box annotations for 37 object categories. ScanNetV2 contains 1.2K training samples, each with axis-aligned bounding box labels belonging to 18 object categories. We extracted 50K points for each of the 6.2K samples, using only the xyz coordinates of each point to construct the pretraining dataset.
Table 1: Classification accuracy on real-scanned (ScanObjectNN) and synthetic (ModelNet40) point clouds. In ScanObjectNN, we report the overall accuracy (%) on three variants. In ModelNet40, we report the overall accuracy (%) for both without and with voting. "#Params" represents the model’s parameter count.
Method Reference #Params (M) ScanObjectNN ModelNet40
OBJ-BG OBJ-ONLY PB-T50-RS w/o Vote w/ Vote
Supervised Learning Only
PointNe (Qi et al., 2017a) CVPR 2017 3.5 73.3 79.2 68 89.2 -
PointNet++ (Qi et al., 2017b) NeurIPS 2017 1.7 82.3 84.3 77.9 90.7 -
DGCNN (Wang et al., 2019) TOG 2019 1.8 82.8 86.2 78.1 92.9 -
PointMLP (Ma et al., 2022) ICLR 2022 12.6 - - 85.2 94.1 94.5
SFR (Zha et al., 2023a) ICASSP 2023 - - - 87.8 93.9 -
P2P-HorNet (Wang et al., 2022) NeurIPS 2022 195.8 - - 89.3 94.0 -
Single Modal Self-Supervised Learning
Point-BERT (Yu et al., 2022) CVPR 2022 22.1 87.43 88.12 83.07 92.7 93.2
MaskPoint (Liu et al., 2022) ECCV 2022 22.1 89.30 88.10 84.30 - 93.8
Point-MAE (Pang et al., 2022) ECCV 2022 22.1 90.02 88.29 85.18 93.2 93.8
Point-M2AE (Zhang et al., 2022) NeurIPS 2022 15.3 91.22 88.81 86.43 93.4 94.0
PointGPT-S (Chen et al., 2024) NeurIPS 2023 29.2 91.63 90.02 86.88 - 94.0
PointDif (Zheng et al., 2024) CVPR 2024 - 93.29 91.91 87.61 - -
MaskFeat3D (Yan et al., 2024) ICLR 2024 15.3 93.20 91.50 88.40 - 94.0
PointMamba (Liang et al., 2024) NeurIPS 2024 12.3 94.32 92.60 89.31 93.6 -
PointHDMAE (Ours) Ours 22.8 95.18 93.12 90.01 93.7 94.2
Multimodal Self-Supervised Learning
Joint-MAE (Guo et al., 2023) IJCAI 2023 - 90.94 88.86 86.07 - 94.0
ACT (Dong et al., 2023) ICLR 2023 22.1 93.29 91.91 88.21 93.2 93.7
TAP+PointMLP (Wang et al., 2023) ICCV 2023 12.6 - - 88.50 94.0 -
I2P-MAE (Zhang et al., 2023) CVPR 2023 15.3 94.15 91.57 90.11 93.7 94.1
Recon (Qi et al., 2023) ICML 2023 44.3 95.18 93.29 90.63 94.1 94.5
Pre-training. During the pretraining phase, we input all 50K×3 point clouds into the scene encoder of PointHDMAE to extract scene-level features. Simultaneously, we randomly select 32 point blocks from the scene point cloud, with each block containing 2K local points, and input these into the 32 object encoders with shared parameters. We use the AdamW (Kingma & Ba, 2014) optimizer with a base learning rate of 5e-4 and a weight decay of 0.1. Simple rotation is applied as data augmentation to both the scene point cloud and each selected object point cloud. For the object point cloud masks, we set the mask ratio to 60% following previous work (Pang et al., 2022; Dong et al., 2023). We train the model from scratch for 200 epochs using 8 A100 GPUs. Furthermore, we can also leverage pre-trained object point cloud models on the ShapeNet55 (Chang et al., 2015) dataset to initialize our object-level models.
4.2 Fine-tuning on Downstream Tasks
4.2.1 Object Point Cloud Classification
We first evaluate the performance of our model on object point cloud classification tasks using the object encoder of PointHDMAE. We conduct point cloud classification on two of the most commonly used object point cloud datasets: ScanObjectNN (Uy et al., 2019) and ModelNet40 (Wu et al., 2015). ScanObjectNN contains 15K real scanned point clouds, each with various backgrounds, occlusions, and noise, which effectively assesses the model’s robustness. ModelNet40 includes 12K synthetic point clouds belonging to 40 different categories, with each point cloud being complete and clean, providing a better representation of 3D object shapes.
Following previous work (Dong et al., 2023; Liang et al., 2024), we use 2K points as input for ScanObjectNN, apply simple rotation for data augmentation, and report classification accuracy without using voting. For ModelNet40, we use 1K points as input, apply scale and translate data augmentation, and report classification accuracy both without voting and with the standard voting mechanism.
As presented in Table 1, firstly, compared to the recent state-of-the-art method PointMamba, our approach surpasses it by 0.86%, 0.52%, and 0.70% on the three variants of ScanObjectNN, respectively. This improvement is significant, given that the same downstream task settings were used. Secondly, our method surpasses the majority of multimodal pre-trained models, ranking just the same with the leading Recon (Qi et al., 2023). This is still highly competitive as Recon (Qi et al., 2023) benefits from the supplementary knowledge of image, and language modalities, while also requiring significantly more parameters than our method. Our PointHDMAE achieves leading performance indicating that using randomly selected point blocks from the scene, despite lacking explicit real-world significance, can still be effectively used for mask reconstruction pre-training. This effectiveness arises because mask reconstruction primarily focuses on learning general representations through the reconstruction of the original shapes, without requiring a specific understanding of the shapes’ concrete meanings.
4.2.2 Scene Point Cloud Detection
We further fine-tune the pre-trained PointHDMAE on scene-level object detection tasks. At this stage, we primarily rely on the scene encoder to process scene-level inputs. Simultaneously, we randomly select 32 point blocks from the scene point cloud and use the 32 shared object encoders to handle these local point blocks. During the decoding phase, we integrate the results from the scene encoder with random queries from the scene, helping the scene encoder to better understand scene details. We use ScanNetV2 (Dai et al., 2017), to evaluate our model’s scene understanding capabilities.
Table 4.2.2 shows our detection results, our PointHDMAE model has shown significant improvements compared with other models. For example, compared to the previous state-of-the-art domain-specific pretraining model PointDif (Zheng et al., 2024), our method achieves a 6.2% improvement on $AP_{50}$. This substantial improvement is mainly attributed to our PointHDMAE using multiple object encoders to assist the scene encoder in analyzing the overall scene. This approach enables the scene model to focus on more local details, thereby enhancing scene understanding.
Table 2: Object detection results on ScanNetV2. We adopt the average precision with 3D IoU thresholds of 0.25 ($AP_{25}$) and 0.5 ($AP_{50}$) for the evaluation metrics.
Methods $AP_{25}$ $AP_{50}$ Supervised Learning Only VoteNet (Qi et al., 2019) 58.6 33.5 3DETR (Misra et al., 2021) 62.1 37.9 Single Modal Self-Supervised Learning PointContrast (Xie et al., 2020b) 58.5 38.0 STRL (Huang et al., 2021) 59.5 38.4 DepthContrast (Zhang et al., 2021) 61.3 - Point-BERT (Yu et al., 2022) 61.0 38.3 MaskPoint (Liu et al., 2022) 64.2 42.1 PointDif (Zheng et al., 2024) - 43.7 PointHDMAE (Ours) 66.8 49.9 Multimodal Self-Supervised Learning PiMAE (Chen et al., 2023) 62.6 39.4 ACT (Dong et al., 2023) 63.8 42.1 TAP (Wang et al., 2023) 62.6 39.4
Table 3: Part segmentation results on the ShapeNetPart. The mean IoU across all categories, i.e., $\mathrm{mIoU}_{c}$ (%), and the mean IoU across all instances, i.e., $\mathrm{mIoU}_{I}$ (%) are reported.
Methods $\mathrm{mIoU}_{c}$ $\mathrm{mIoU}_{I}$ Supervised Learning Only PointNet++ (Qi et al., 2017b) 81.9 85.1 PointMLP (Ma et al., 2022) 84.6 86.1 Single-Modal Self-Supervised Learning PointContrast (Xie et al., 2020b) - 85.1 CrossPoint (Afham et al., 2022) - 85.5 Point-BERT (Yu et al., 2022) 84.1 85.6 MaskPoint (Liu et al., 2022) 84.4 86.0 Point-MAE (Pang et al., 2022) 84.2 86.1 PointGPT-S (Chen et al., 2024) 84.1 86.2 PointMamba (Liang et al., 2024) 84.4 86.2 PointHDMAE (Ours) 85.0 86.3 Multimodal Self-Supervised Learning ACT (Dong et al., 2023) 84.7 86.1 Joint-MAE (Guo et al., 2023) 85.4 86.3 Recon (Qi et al., 2023) 84.8 86.4
4.2.3 Object Point Cloud Part Segmentation
We assess the performance of our PointHDMAE in part segmentation using the ShapeNetPart dataset (Chang et al., 2015), comprising 16,881 samples across 16 categories. We utilize the same segmentation setting after the pre-trained encoder as in previous works Pang et al. (2022); Zhang et al. (2022) for fair comparison. The experimental results, displayed in Table 4.2.2, demonstrate that our model exhibits competitive performance in tasks such as part segmentation, which demands a more fine-grained understanding of point clouds.
4.2.4 Object Point Cloud Completion
Previous pretraining models have not explored the effects on low-level tasks. We first specifically design task heads for downstream low-level point cloud completion tasks. Then, we fine-tune our PointHDMAE model on point cloud completion on the classic point cloud completion dataset PCN (Yuan et al., 2018) and ShapeNet-55 (Chang et al., 2015). The PCN dataset is created from the ShapeNet (Chang et al., 2015) dataset, including eight categories with a total of 30974 CAD models. Compared to the PCN dataset, ShapeNet-55 includes 55 different categories of 3D models. Following previous practices (Yu et al., 2021; Li et al., 2023), we used 41,952 models for training and 10,518 models for testing. For each object, we randomly sampled 8,192 points from the surface to obtain the point cloud. We also divide the test samples into three difficulty degrees, simple, moderate, and hard in our experiments and we report the performance for each method in simple, moderate, and hard to show the ability of each network to deal with tasks at difficulty levels.
We followed the data processing methods established in previous works (Yu et al., 2021; Li et al., 2023) and report the average $l_{1}$ Chamfer Distance (Fan et al., 2017) (CD-Avg) across all 8 object categories in the PCN dataset, and the average Chamfer Distance for the 55 categories under the simple (CD-S), moderate (CD-M), and hard (CD-H) settings in the ShapeNet-55 dataset, along with the overall average of these three metrics (CD-Avg). As shown in table 4, our model achieves state-of-the-art results across various settings in both datasets, demonstrating that our pre-trained model can better handle diverse scenarios, such as different viewpoints, object categories, incomplete patterns, and varying levels of incompleteness, even in lower-level completion tasks.
Table 4: Quantitative comparison of point cloud completion task on PCN. Point resolutions for the output and ground-truth are 16384. For Chamfer Distance, lower is better.
Methods Reference PCN ShapeNet-55
CD-Avg CD-S CD-M CD-H CD-Avg
FoldingNet (Yang et al., 2018) CVPR 2018 14.31 2.67 2.66 4.05 3.12
PCN (Yuan et al., 2018) 3DV 2018 9.64 1.94 1.96 4.08 2.66
GRNet (Xie et al., 2020a) ECCV 2020 8.83 1.35 1.71 2.85 1.97
PoinTr (Yu et al., 2021) ICCV 2021 8.38 0.58 0.88 1.79 1.09
LAKeNet (Tang et al., 2022) CVPR 2022 7.23 - - - 0.89
SnowFlakeNet (Xiang et al., 2021) ICCV 2021 7.21 0.70 1.06 1.96 1.24
ProxyFormer (Li et al., 2023) CVPR 2023 6.77 0.49 0.75 1.55 0.93
SeedFormer (Zhou et al., 2022) ECCV 2022 6.74 0.50 0.77 1.49 0.92
PointHDMAE Ours 6.54 0.51 0.70 1.24 0.81
4.2.5 Scene Semantic Segmentation
We have conducted experiments on scene-level semantic segmentation tasks to assess the performance of PointHDMAE in classifying each point in a scene into semantic categories. We validated our model using the indoor S3DIS (Armeni et al., 2016) dataset for semantic segmentation tasks. Specifically, we tested the model on Area 5 while training on other areas and report the mean IoU (mIoU) and mean Accuracy (mAcc). To ensure a fair comparison, we used the same codebase based on the PointNext (Qian et al., 2022a) baseline and employed identical decoders and semantic segmentation heads. We acknowledge that there are several outstanding works in semantic segmentation, such as PointTransformerV3 (Wu et al., 2024), that achieve performance far exceeding that of PointNeXt. However, these advanced models often require more complex inputs, such as color and normal information. Our focus is on segmentation using only the most basic xyz coordinate information. Therefore, we chose PointNeXt as the baseline codebase for our pretraining model and ensure a fair comparison with other pretraining models.
The experimental results are shown in the table 4.2.5. Compared to training the PointNeXt model from scratch, our method improves the mIoU score by 2.3%. It also shows significant improvements over other pretraining models, such as Point-MAE (Pang et al., 2022) and PointDif (Zheng et al., 2024). This enhancement is largely due to our block-to-scene pretraining, which equips the model with strong scene understanding capabilities and further demonstrates the generalizability of our approach.
Table 5: Semantic segmentation results on S3DIS Area 5.
Methods mIoU mAcc PointCNN (Li et al., 2018) 57.3 63.9 Pix4Point (Qian et al., 2022b) 69.6 75.2 PointNeXt (Qian et al., 2022a) 68.5 75.1 Point-BERT (Yu et al., 2022) 68.9 76.1 MaskPoint (Liu et al., 2022) 68.6 74.2 Point-MAE (Pang et al., 2022) 68.4 76.2 PointDif (Zheng et al., 2024) 70.0 77.1 PointHDMAE (Ours) 70.8 77.6
Table 6: Head tuning of object-level classification and scene-level detection.
Object-Level Classification Methods OBJ-BG OBJ-ONLY Point-BERT (Yu et al., 2022) 88.81 88.3 Point-MAE (Pang et al., 2022) 89.50 88.98 PointHDMAE (Ours) 90.71 88.47 Scene-Level Detection Methods $AP_{25}$ $AP_{50}$ Point-BERT (Yu et al., 2022) 60.13 38.5 Point-MAE (Pang et al., 2022) 60.82 40.4 PointHDMAE (Ours) 62.44 41.8
4.2.6 Head Tuning on Downstream Tasks
We further demonstrate the generalization capability of our pre-trained model by performing only head tuning on downstream tasks. We keep the pre-trained feature extractor fixed and only train the task-specific heads (including the classification head and detection head). At the same time, all models use the same downstream task setting as described in 4.2. These experiments are designed to better evaluate the performance of the proposed PointHDMAE and to facilitate a more direct comparison with the baselines. The results of these experiments are summarized in the tables 4.2.5. As shown, our proposed Point-HDMAE still demonstrates superior performance compared to the baselines, highlighting its robust representational capabilities even without fine-tuning the entire model.
4.3 Ablation Study
4.3.1 The Impact of Stop Gradients.
In our implementation, stopping gradients is crucial. This approach allows different encoders within the model to learn scene-level and object-level representations independently during block-to-scene pretraining. By ensuring that each encoder focuses solely on its specific task, we enhance the model’s generalization capability. Consequently, maintaining the accuracy and independence of each encoder’s learning objective during training is essential.
Since the position regression objective and the masked reconstruction objective are two distinct tasks in our pretraining process, failing to decouple the learning processes of the different encoders could lead to catastrophic forgetting. For instance, without gradient stopping, gradients from the object-level reconstruction tasks could backpropagate into the scene’s encoder, interfering with its ability to learn scene-level knowledge. By applying gradient stopping, we effectively prevent this interference, ensuring that each encoder remains focused on its specific task and thereby avoiding catastrophic forgetting.
To further validate this issue, we conducted experiments where we removed the Stop Gradients operation from our pipeline and retrained the PointHDMAE model. We then assessed its performance on both scene-level point cloud detection tasks and object-level classification tasks. This comparison allowed us to observe the impact of gradient stopping on the model’s ability to effectively learn and generalize across different tasks. As shown in table 4.3.1, after removing the Stop Gradients operation, there was a 8.6% decrease in AP50 for the scene-level detection tasks, clearly indicating a deterioration in the representation capabilities of the scene encoder. Similarly, there was a noticeable decline in performance across classification tasks. These results collectively demonstrate that decoupling the representation learning of different encoders is essential to avoid catastrophic forgetting.
Table 7: The impact of stop gradients.
Object-Level Classification OBJ-BG OBJ-ONLY w/o Stop Gradients 93.29 92.43 w/ Stop Gradients 95.18 93.12 Scene-Level Detection $AP_{25}$ $AP_{50}$ w/o Stop Gradients 63.1 41.3 w/ Stop Gradients 66.8 49.9
Table 8: The impact of the number of point blocks.
Number of blocks FLOPs(G) PB-T50-RS 1 22 88.78 8 59 89.10 16 102 89.52 32 187 90.01 64 359 90.12
4.3.2 The Impact of the Number of Point Blocks.
The number of randomly selected point blocks during the pretraining phase has a significant impact on the representation learning of the object encoder. Each point block serves as a sample to train the object encoder. The more point blocks selected from a scene, the richer the knowledge the object encoder can learn. However, this also leads to a significant increase in computational complexity. Therefore, we need to select an appropriate number to achieve a balance between efficiency and performance.
We selected 1, 8, 16, 32, and 64 blocks respectively and pre-trained these models from scratch. We reported the computational floating-point operations (FLOPs) required for pre-training and the performance of the resulting object models. The trained models were fine-tuned on the PB-RS-T50 variant of ScanObjectNN, using its split to evaluate model performance. As illustrated in the table 4.3.1, our model’s computational complexity significantly increases with the number of blocks, while the performance of the model gradually improves and eventually levels off. To achieve a balance between performance and efficiency, we chose 32 blocks for our experiments.
4.4 Conclusion
In this paper, we first propose a point cloud hybrid-domain masked autoencoders model to address the generalization limitations of existing domain-specific models. Our hybrid model selectively activates the object encoder to handle object domain point clouds specifically and leverages these object encoders to assist the scene encoder in processing scene domain point clouds. Then, we propose a block-to-scene pre-training strategy to train our PointHDMAE model. This strategy involves joint training through random object mask reconstruction and position regression within the scene, enabling our domain-specific encoder models to learn general representations relevant to their respective domains. Finally, we demonstrated the generalization and superiority of our model through extensive experiments across various datasets and tasks from different domains.

### v3

@@
-Block-to-Scene Pre-training for Point Cloud Hybrid-Domain Masked Autoencoders
+Point Cloud Mixture-of-Domain-Experts Model for 3D Self-supervised Learning
 Abstract
-Point clouds, as a primary representation of 3D data, can be categorized into scene domain point clouds and object domain point clouds based on the modeled content.
-Masked autoencoders (MAE) have become the mainstream paradigm in point clouds self-supervised learning.
-However, existing MAE-based methods are domain-specific, limiting the model’s generalization.
-In this paper, we propose to pre-train a general Point cloud Hybrid-Domain Masked AutoEncoder (PointHDMAE) via a block-to-scene pre-training strategy.
-We first propose a hybrid-domain masked autoencoder consisting of an encoder and decoder belonging to the scene domain and object domain, respectively.
-The object domain encoder specializes in handling object point clouds and multiple shared object encoders assist the scene domain encoder in analyzing the scene point clouds.
-Furthermore, we propose a block-to-scene strategy to pre-train our hybrid-domain model.
-Specifically, we first randomly select point blocks within a scene and apply a set of transformations to convert each point block coordinates from the scene space to the object space.
-Then, we employ an object-level mask and reconstruction pipeline to recover the masked points of each block, enabling the object encoder to learn a universal object representation.
-Finally, we introduce a scene-level block position regression pipeline, which utilizes the blocks’ features in the object space to regress these blocks’ initial positions within the scene space, facilitating the learning of scene representations.
-Extensive experiments across different datasets and tasks demonstrate the generalization and superiority of our hybrid-domain model.
-The code will be released.
+Point clouds, as a primary representation of 3D data, can be categorized into scene domain point clouds and object domain point clouds.
+Point cloud self-supervised learning (SSL) has become a mainstream paradigm for learning 3D representations.
+However, existing point cloud SSL primarily focuses on learning domain-specific 3D representations within a single domain, neglecting the complementary nature of cross-domain knowledge, which limits the learning of 3D representations.
+In this paper, we propose to learn a comprehensive Point cloud Mixture-of-Domain-Experts model (Point-MoDE) via a block-to-scene pre-training strategy.
+Specifically, We first propose a mixture-of-domain-expert model consisting of scene domain experts and multiple shared object domain experts.
+Furthermore, we propose a block-to-scene pretraining strategy, which leverages the features of point blocks in the object domain to regress their initial positions in the scene domain through object-level block mask reconstruction and scene-level block position regression.
+By integrating the complementary knowledge between object and scene, this strategy simultaneously facilitates the learning of both object-domain and scene-domain representations, leading to a more comprehensive 3D representation.
+Extensive experiments in downstream tasks demonstrate the superiority of our model.
 1 Introduction
-With the rapid development of 3D scanning technology, 3D point clouds have become the mainstream representation for 3D objects due to their ease of acquisition, explicit representation, and efficient storage.
-Point clouds can be categorized into scene domain point clouds (Dai et al., 2017; Song et al., 2015; Armeni et al., 2016; Zheng et al., 2020; Sun et al., 2020) and object domain point clouds (Wu et al., 2015; Chang et al., 2015; Uy et al., 2019; Deitke et al., 2024; Yu et al., 2023) based on the modeling object.
+Recently, with the rapid development of 3D scanning technology, 3D point clouds have become the mainstream representation for 3D objects due to their ease of acquisition, explicit representation, and efficient storage.
+Point clouds can be categorized into scene domain point clouds Dai et al.
+(2017); Song et al.
+(2015); Armeni et al.
+(2016) and object domain point clouds Wu et al.
+(2015); Chang et al.
+(2015); Uy et al.
+(2019) based on the modeling object.
 As shown in Figure 2 (a), object domain point clouds describe specific objects or entities, such as an airplane, with relatively fewer points.
@@
 Due to the significant disparity in point count and the elements being described, a substantial domain gap exists in these two types of point clouds.
-Figure 1: The handling approach for point clouds from different domains in domain-specific models (a) compared to our hybrid-domain model (b), and the generalization experiments of the domain-specific Point-MAE model (c).
-Recently, point cloud masked autoencoders (Yu et al., 2022; Pang et al., 2022; Zhang et al., 2022; Dong et al., 2023; Zha et al., 2024a), pre-trained on massive point cloud data, have become the mainstream paradigm in point cloud self-supervised learning and have been widely applied to various point cloud tasks.
-It is inspired by masked image modeling (Bao et al., 2021; He et al., 2022; Xie et al., 2022), using the unmasked portions to predict the geometric coordinates or semantic features of the masked parts, thereby enabling the model to learn universal 3D representations.
+Figure 1: Comparison of (a) the traditional domain-specific model and (b) our mixture-of-domain-experts model for processing point clouds from different domains.
+Point cloud self-supervised learning (SSL) Yu et al.
+(2022); Pang et al.
+(2022); Zhang et al.
+(2022); Dong et al.
+(2023), pre-trained on massive point cloud data, have become the mainstream paradigm for learning 3D representations and have been widely transferred to various point cloud tasks Zhou et al.
+(2024); Zha et al.
+(2025).
 Despite the significant success, most of these methods are domain-specific due to the notable domain gap between object-domain and scene-domain point clouds.
-As shown in Figure 1 (a), these methods (Yu et al., 2022; He et al., 2022; Zhang et al., 2022; Zha et al., 2023b; Zhou et al., 2024) use scene-level models for scene tasks and object-level models for object tasks, thereby limiting their generalizability.
-For example, Point-MAE (Pang et al., 2022), which is pre-trained on ShapeNet (Chang et al., 2015), primarily performs object point cloud tasks.
-For scene point clouds, it requires re-pretraining on scene-level datasets like ScanNet (Dai et al., 2017) to adapt to the scene domain.
-As shown in Figure 1 (c), directly transferring an object domain pre-trained model (e.g.Point-MAE (Pang et al., 2022)) to scene tasks results in a significant performance drop.
+As shown in Figure 2 (a), these methods Yu et al.
+(2022); He et al.
+(2022); Zhang et al.
+(2022) use scene-level data to learn scene model and object-level data for object model.
+This domain-independent learning approach not only fails to effectively integrate knowledge across different domains, but also struggles to perform well on data from multiple domains simultaneously.
+For example, Point-MAE Pang et al.
+(2022), which is pre-trained on ShapeNet Chang et al.
+(2015), primarily performs object point cloud tasks.
+For scene point clouds, it requires re-pretraining on scene-level datasets like ScanNet Dai et al.
+(2017) to adapt to the scene domain.
+As shown in Figure 2 (c), directly transferring an object domain Point-MAE to scene tasks results in a significant performance drop.
 Similarly, models pre-trained on the scene domain also exhibit performance declines when transferred to object task.
-Pre-training a general point cloud model is our persistent pursuit; however, it is highly challenging for two main reasons.
+Figure 2: Transferability of the domain-specific Point-MAE model.
+Transferring a Point-MAE pre-trained on the scene to an object task results in a clear performance drop, and the same occurs for models in the object domain..
+Learning a more comprehensive 3D representation by simultaneously integrating knowledge from both the object domain and the scene domain is a reasonable solution.
+However, it is highly challenging for two main reasons.
 Firstly, the input data is inconsistent.
-Scene-level point clouds, such as ScanNet (Dai et al., 2017), typically consist of 50k points, while object-level point clouds like ModelNet (Wu et al., 2015) typically consist of 1k points.
+Scene-level point clouds, such as ScanNet Dai et al.
+(2017), typically consist of 50k points, while object-level point clouds like ModelNet Wu et al.
+(2015) typically consist of 1k points.
 The disparity in point count between the two types is significant, making it difficult to process both types of data simultaneously using a single model.
@@
 On the contrary, object point clouds generally involve classification tasks, which tend to prioritize understanding global geometry.
-To address the aforementioned challenges, we propose a block-to-scene pretraining strategy to pre-train a Point cloud Hybrid-Domain Masked Auto-Encoder (PointHDMAE).
-We address the challenge of inconsistent input data by using domain-specific encoders to process data from their respective domains.
+To address the aforementioned challenges, we propose a block-to-scene pretraining strategy to pre-train a Point cloud Mixture-of-Domain-Experts model (Point-MoDE).
+We address the challenge of inconsistent input data by using domain-specific experts to process data from their respective domains.
 Additionally, we finetune the pre-trained model to address the inconsistency of task emphasis.
-Specifically, as shown in Figure 2, we first design a point cloud hybrid-domain architecture that consists of an encoder and decoder belonging to the scene domain and object domain, respectively.
-In the fine-tuning phase, as shown in Figure 1 (b), for object domain data, our model selectively activates the object-domain encoder for analysis.
-However, in the case of scene point clouds, we activate multiple shared object encoders to assist the scene encoder in analyzing scene domain data collaboratively.
-Furthermore, we propose a block-to-scene pre-training strategy that couples masked reconstruction and position regression tasks of random object blocks within a scene for self-supervised learning, enabling us to train encoders for different domains simultaneously.
+Specifically, as shown in Figure 2 (b), we first design a point cloud mixture-of-domain-experts model that consists of a scene domain expert and multiple shared object domain experts.
+In the fine-tuning phase, for object domain data, our model selectively activates the object-domain expert for analysis.
+However, in the case of scene point clouds, we activate multiple shared object expert encoders to assist the scene expert in analyzing scene domain data collaboratively.
+Furthermore, we propose a block-to-scene pre-training strategy that couples masked reconstruction and position regression tasks of random object blocks within a scene for self-supervised training, enabling us to train experts for different domains simultaneously.
 Specifically, we first randomly select point blocks within a scene and apply a set of transformations to convert each point block coordinates from the scene space to the object space.
-Then, within the object domain, we use a mask and reconstruction pipeline to recover the masked points of each block, enabling it to learn universal object representations.
-Finally, we introduce a scene-level block position regression pipeline, which utilizes the blocks’ features in the object space to regress these blocks’ initial positions within the scene space, enabling the scene encoder to learn scene representations with the assistance of the object encoders.
-By block-to-scene pretraining, our model can simultaneously learn powerful object-level and scene-level representations and exhibit superior transferability.
+Then, within the object expert, we use a mask and reconstruction pipeline to recover the masked points of each block, enabling it to learn universal object representations.
+Finally, we introduce a scene-expert-based block position regression pipeline, which utilizes the blocks’ features in the object space to regress these blocks’ initial positions within the scene space, enabling the scene expert to learn scene representations with the assistance of the object experts.
+By block-to-scene pretraining, our Point-MoDE can simultaneously learn powerful object-level and scene-level representations and exhibit superior transferability.
 Our model can be fine-tuned directly on downstream tasks such as object point cloud classification, segmentation, completion, and scene point cloud detection without the need for any additional domain adaptation training.
-Our main contributions can be summarized as follows: (1) We propose a point cloud hybrid-domain masked autoencoders to address the generalization limitations of existing domain-specific MAE-based models.
-(2) We propose a block-to-scene pretraining strategy, a joint pre-training strategy that reconstructs and regresses random object blocks within a scene.
-(3) Extensive experiments across different datasets and tasks demonstrate the generalization and superiority of our model.
+The main contributions can be summarized as follows:
+  • We propose a mixture-of-domain-experts model, integrating both scene domain and object domain experts to effectively process point clouds from different domains.
+  • We introduce a novel block-to-scene pretraining strategy that combines masked reconstruction and position regression tasks, enabling the simultaneous learning of object-level and scene-level representations.
+  • Extensive experiments across different datasets and tasks demonstrate the superiority and transferability of our model.
+Figure 3: The architecture of our point cloud mixture-of-domain-experts model and the pipeline for block-to-scene pre-training.
+The left side illustrates the scene-level block position regression, while the right side shows the object-level block masked reconstruction.
 2 Related Work
-2.1 Self-supervised Learning for Point Cloud.
-Self-supervised learning, which enables the learning of general representations from large amounts of unlabeled data, has been widely applied in fields such as language (Semnani et al., 2019; Brown et al., 2020; Achiam et al., 2023) and image (Bao et al., 2021; Chen et al., 2020b; a; He et al., 2022).
+2.1 Self-supervised Learning for Point Cloud
+Self-supervised learning, which enables the learning of general representations from large amounts of unlabeled data, has been widely applied in fields such as language Semnani et al.
+(2019); Brown et al.
+(2020); Achiam et al.
+(2023) and image Bao et al.
+(2021); Chen et al.
+(2020b, a); He et al.
+(2022).
 Inspired by the success of visual pretraining, numerous point cloud pretraining methods have also been proposed.
-Based on the pretext tasks, they can be categorized into contrastive learning paradigms (Oord et al., 2018; Tian et al., 2020) and masked reconstruction paradigms (Bao et al., 2021; He et al., 2022).
-PointContrast (Xie et al., 2020b), CrossPoint (Afham et al., 2022), and DepthContrast (Zhang et al., 2021) construct positive and negative sample pairs using various methods and employ contrastive learning techniques to learn 3D representations.
-Point-BERT (Yu et al., 2022) was the first to propose learning universal 3D representations using the paradigm of masked reconstruction.
-Subsequently, numerous explorations (Pang et al., 2022; Zhang et al., 2022; Zha et al., 2024a; b) have improved masked reconstruction from various perspectives.
-Point-MAE (Pang et al., 2022) and PointFEMAE (Zha et al., 2024a) introduced the masked autoencoder for reconstruction, and PointGPT (Chen et al., 2024) proposed pretraining using an autoregressive approach.
+Based on the pretext tasks, they can be categorized into contrastive learning paradigms Oord et al.
+(2018); Tian et al.
+(2020) and masked reconstruction paradigms Bao et al.
+(2021); He et al.
+(2022).
+PointContrast Xie et al.
+(2020), CrossPoint Afham et al.
+(2022), and DepthContrast Zhang et al.
+(2021) construct positive and negative sample pairs using various methods and employ contrastive learning techniques to learn 3D representations.
+Point-BERT Yu et al.
+(2022) was the first to propose learning universal 3D representations using the paradigm of masked reconstruction.
+Subsequently, numerous explorations have improved masked reconstruction from various perspectives.
+Point-MAE Pang et al.
+(2022) and Point-M2AE Zhang et al.
+(2022) introduced the masked autoencoder for reconstruction, and PointGPT Chen et al.
+(2024) proposed pretraining using an autoregressive approach.
 To address the limited amount of point cloud during pretraining, many approaches integrate multimodal knowledge to aid in learning point cloud features.
-ACT (Dong et al., 2023) leverages a pre-trained image model to assist in point cloud reconstruction, while I2P-MAE (Zhang et al., 2023) employs an image-guided masking strategy.
-PiMAE (Chen et al., 2023) proposes to address the challenges of multi-modal interaction between point cloud and RGB image data through mask alignment, a two-branch MAE pipeline, and a cross-modal reconstruction module.
-In this paper, we propose a block-to-scene pretraining strategy that combines masked reconstruction and position regression in a joint self-supervised learning method to enhance the model’s generalizability.
+ACT Dong et al.
+(2023) leverages a pre-trained image model to assist in point cloud reconstruction, while I2P-MAE Zhang et al.
+(2023) employs an image-guided masking strategy.
+Despite the significant success, the above methods are domain-specific.
+In this paper, we propose training a mixture-of-domain-experts model to integrate knowledge from different domains for learning comprehensive 3D representations.
 3 Methodology
-In this section, we provide a detailed explanation of how to use our block-to-scene pretraining strategy to train our point cloud hybrid-domain masked autoencoder.
-3.1 Point Cloud Hybrid-Domain Masked Autoencoder (PointHDMAE)
-The overall architecture of our PointHDMAE is shown in Figure 2, it is composed of four main components: a scene encoder, a scene decoder, a shared object encoder, and a shared object decoder.
+In this section, we provide a detailed explanation of how to use our block-to-scene pretraining strategy to train our point cloud mixture-of-domain-experts model.
+3.1 Point Cloud Mixture-of-Domain-Experts
+The overall architecture of our Point-MoDE is shown in Figure 3, it is composed of four main components: a scene expert encoder, a scene expert decoder, a shared object expert encoder, and a shared object expert decoder.
 It is primarily used for two main task pipelines: object point cloud processing and scene point cloud processing.
-Our PointHDMAE is a hybrid model.
+Our Point-MoDE is a hybrid model.
 Due to significant differences across domains of point clouds and tasks, we selectively activate different sub-modules for various downstream data and tasks.
-For instance, in tasks such as object point cloud classification and object part segmentation, we selectively activate our object encoder according to specific tasks.
+For instance, in tasks such as object point cloud classification and object part segmentation, we selectively activate our object expert encoder according to specific tasks.
 For scene point cloud detection tasks, we activate all encoders.
-This collaborative approach is primarily adopted because utilizing the object encoder for analyzing local point blocks within the scene contributes to the scene encoder’s comprehension of scene intricacies.
+This collaborative approach is primarily adopted because utilizing the object expert encoder for analyzing local point blocks within the scene contributes to the scene encoder’s comprehension of scene intricacies.
+In the scene expert model, we adopt a standard Transformer as our scene expert, comprising 3 layers of Transformer-based encoders and 8 layers of Transformer-based decoders with a PointNet-based Qi et al.
+(2017a) token embedding layer.
+In our object expert model, we directly leverage existing pre-trained object models based on mask-based point modeling, such as Point-BERT Yu et al.
+(2022), Point-MAE Pang et al.
+(2022), PointGPT Chen et al.
+(2024), etc.
+Using these models pre-trained on the object domain allows us to incorporate object-level priors into our pipeline while saving training resources and enabling better extension of existing models.
+This characteristic makes our object experts flexible, allowing us to replace them with different backbones.
 3.2 Block-to-Scene Pretraining
@@
 Below, we provide a detailed illustration of the specific implementation of each component.
-Figure 2: The architecture of our point cloud hybrid-domain masked autoencoder and the pipeline for block-to-scene pre-training.
-The left side illustrates the scene-level block position regression, while the right side shows the object-level block masked reconstruction.
 3.2.1 Random Point Block Generation
@@
 We generate the ground truth position of each random point block in the scene, which will be used to constrain the predicted position regressed by the final scene decoder.
-Inspired by the detection (Carion et al., 2020; Misra et al., 2021; Dai et al., 2021) task, we use the 3D bounding box of each random point block as its ground truth position.
+Inspired by the detection Carion et al.
+(2020); Misra et al.
+(2021); Dai et al.
+(2021) task, we use the 3D bounding box of each random point block as its ground truth position.
 By computing the mean of all points in each dimension of the entire point block, the coordinates of the center point are obtained.
@@
 Subsequently, the center point coordinates and half-lengths in each dimension (x, y, and z) are concatenated to form the bounding box.
-Finally, standard procedures (Misra et al., 2021) are applied to compute bounding box parameters $\bm{B_{b}}$ such as size and corners for each bounding box.
+Finally, standard procedures Misra et al.
+(2021) are applied to compute bounding box parameters $\bm{B_{b}}$ such as size and corners for each bounding box.
 3.2.2 Object-level Block Masked Reconstruction
+We first initialize our object expert model using the Point-MAE (or Point-BERT, PointGPT, etc.) model pre-trained on ShapeNet Chang et al.
+(2015).
+Since these models are open-source, we directly use the pre-trained weights provided by the official repository.
+This initialization allows us to incorporate object-level knowledge priors into our pipeline without the need for additional object data or specialized training.
 Object point block generation.
-We treat each randomly selected point block in $\bm{B}$ as an object point block and transform its coordinates from the scene space to the object space for processing by the object autoencoder.
+We treat each randomly selected point block in $\bm{B}$ as an object point block and transform its coordinates from the scene space to the object space for processing by the object expert.
 Specifically, we apply a simple set of transformation functions to each point block.
@@
 Then, we normalize these coordinates to the range [-1, 1].
-Finally, after applying a random rotation transformation to each point block, we obtain all point blocks $\bm{B_{o}=\{B_{o}^{1},...,B_{o}^{K_{o}}\}}$ as input to the object encoder.
+Finally, after applying a random rotation transformation to each point block, we obtain all point blocks $\bm{B_{o}=\{B_{o}^{1},...,B_{o}^{K_{o}}\}}$ as input to the object expert encoder.
 Through these transformation functions, we convert the coordinates of each point block from the scene space to the object space, decoupling the point block object coordinates from the original scene coordinates.
-This enables the object encoder to learn the universal shape features of the point block objects.
+This enables the object expert to learn the universal shape features of the point block objects.
 Object point block masked and reconstruction.
-We use a shared object autoencoder to perform mask-based reconstruction self-supervised learning on all generated object point blocks $\bm{B_{o}}$, enabling our object encoder to learn a general representation of objects.
-We illustrate the entire mask and reconstruction process for the object point blocks using Point-MAE (Pang et al., 2022) as an example.
+We use a shared object autoencoder to perform mask-based reconstruction self-supervised learning on all generated object point blocks $\bm{B_{o}}$, enabling our object expert to learn a general representation of objects.
+We illustrate the entire mask and reconstruction process for the object point blocks using Point-MAE Pang et al.
+(2022) as an example.
 For the $i$-th object point block $\bm{B_{o}^{i}\in\mathbb{R}^{N_{o}\times 3}}$, we use farthest point sampling and the K-nearest neighbors algorithm to divide it into $M_{o}$ point patches.
@@
 By adding them, we obtain the token $\bm{E}_{o}^{0}\in\mathbb{R}^{rM_{o}\times C_{o}}$ for each unmasked patch, where $r$ represents the unmasked ratio, and $C_{o}$ denotes the object feature dimension.
-Finally, we use a shared object encoder to extract object features $\bm{E}_{o}^{n_{o}}\in\mathbb{R}^{rM_{o}\times C_{o}}$, where $n_{o}$ is the number of layers in the scene encoder.
+Finally, we use a shared object expert encoder to extract object features $\bm{E}_{o}^{n_{o}}\in\mathbb{R}^{rM_{o}\times C_{o}}$, where $n_{o}$ is the number of layers in the scene expert.
 In the decoding phase, we concatenate $\bm{E}_{o}^{n_{o}}$ with randomly initialized masked tokens to obtain $\bm{D}_{o}^{0}\in\mathbb{R}^{M_{o}\times C_{o}}$.
@@
 By adding them, we obtain the token $\bm{E}_{s}^{0}\in\mathbb{R}^{M_{s}\times C_{s}}$ for each patch, where $M_{s}$ represents the number of scene patches, and $C_{s}$ denotes the scene feature dimension.
-Finally, we use a scene encoder based on the standard Transformer (Vaswani et al., 2017) architecture to extract scene features $\bm{E}_{s}^{n_{s}}\in\mathbb{R}^{M_{s}\times C_{s}}$, where $n_{s}$ is the number of Transformer layers in the scene encoder.
+Finally, we use a scene encoder based on the standard Transformer Vaswani et al.
+(2017) architecture to extract scene features $\bm{E}_{s}^{n_{s}}\in\mathbb{R}^{M_{s}\times C_{s}}$, where $n_{s}$ is the number of Transformer layers in the scene encoder.
 Scene decoding and position regression.
@@
 After passing through the projection layer, these point block features are transformed into features $\bm{B}_{g}\in\mathbb{R}^{K_{o}\times C_{s}}$ for scene decoding input.
-However, we prevent the gradients of $\bm{B}_{g}$ from propagating backward into the mask reconstruction pipeline during the backpropagation process, thereby mitigating the multi-task interference caused by the scene regression task on the object encoder.
+However, we stop the gradients of $\bm{B}_{g}$ from propagating backward into the mask reconstruction pipeline during the backpropagation process, thereby mitigating the multi-task interference caused by the scene regression task on the object encoder.
 We will provide a detailed explanation of this issue in Section 4.4.2.
@@
 Since the number of point blocks and queries often differ, we replicate $\bm{B}_{g}$ to match all queries.
-We use a Transformer decoder based on self-attention and cross-attention as our scene decoder.
-The input queries $\bm{Q^{0}}$, with the assistance of the encoded features $\bm{E}_{s}^{n_{s}}$, pass through our decoder to obtain the decoded query features $\bm{Q^{m_{s}}}$, where $m_{s}$ is the number of scene decoder.
+We use a Transformer decoder based on self-attention and cross-attention as our scene expert decoder.
+The input queries $\bm{Q^{0}}$, with the assistance of the encoded features $\bm{E}_{s}^{n_{s}}$, pass through our decoder to obtain the decoded query features $\bm{Q^{m_{s}}}$, where $m_{s}$ is the number of scene expert decoder.
 Finally, we use different MLP-based reconstruction heads to predict the 3D bounding box $\bm{B_{b}^{p}}$ of each random point block.
@@
 We use a combination of mask reconstruction loss and point block regression loss to jointly constrain our pre-training process.
-For the reconstruction loss calculation, we follow previous work (Pang et al., 2022; Zhang et al., 2022) and use Chamfer Distance (Fan et al., 2017) (CD) as the loss function.
-For the regression loss calculation, we reference detection (Misra et al., 2021) and use Generalized Intersection over Union (Rezatofighi et al., 2019) (GIoU) as the loss function.
+For the reconstruction loss calculation, we follow previous work Pang et al.
+(2022); Zhang et al.
+(2022) and use Chamfer Distance Fan et al.
+(2017) (CD) as the loss function.
+For the regression loss calculation, we use Generalized Intersection over Union Rezatofighi et al.
+(2019) (GIoU) as the loss function.
 Therefore, our loss function is defined as follows:
@@
 where $\lambda_{1}$ and $\lambda_{2}$ is a weighted combination of reconstruction loss and regression loss.
+Method Reference #Params (M) ScanObjectNN ModelNet40
+OBJ-BG OBJ-ONLY PB-T50-RS
+Supervised Learning Only
+PointNet Qi et al. (2017a) CVPR 2017 3.5 73.3 79.2 68.0 89.2
+PointNet++ Qi et al. (2017b) NeurIPS 2017 1.7 82.3 84.3 77.9 90.7
+SFR Zha et al. (2023a) ICASSP 2023 - - - 87.8 93.9
+PointMLP Ma et al. (2022) ICLR 2022 12.6 - - 85.2 94.5
+Self-Supervised Learning
+Point-BERT Yu et al. (2022) CVPR 2022 22.1 87.43 88.12 83.07 93.2
+Point-MAE Pang et al. (2022) ECCV 2022 22.1 90.02 88.29 85.18 93.8
+Point-M2AE Zhang et al. (2022) NeurIPS 2022 15.3 91.22 88.81 86.43 94.0
+PointGPT-S Chen et al. (2024) NeurIPS 2023 29.2 91.6 90.0 86.9 94.0
+PointDif Zheng et al. (2024) CVPR 2024 - 93.29 91.91 87.61 -
+ACT Dong et al. (2023) ICLR 2023 22.1 93.29 91.91 88.21 93.7
+LCM Zha et al. (2024b) NeurIPS 2024 2.7 94.51 92.75 88.87 94.2
+Point-FEMAE Zha et al. (2024a) AAAI 2024 27.4 95.18 93.29 90.22 94.5
+Point-BERT (baseline) CVPR 2022 22.1 92.48 91.60 87.91 93.2
+Point-MAE (baseline) ECCV 2022 22.1 92.67 92.08 88.27 93.8
+PointGPT-B (baseline) NeurIPS 2023 120.5 95.8 95.2 91.9 94.4
+Point-MoDE w/ Point-BERT Ours 22.1 93.46($\uparrow$ 1.0) 92.25($\uparrow$ 0.7) 88.58($\uparrow$ 0.7) 93.6($\uparrow$ 0.4)
+Point-MoDE w/ Point-MAE Ours 22.1 93.98($\uparrow$ 1.3) 92.77($\uparrow$ 0.7) 89.14($\uparrow$ 0.9) 94.0($\uparrow$ 0.2)
+Point-MoDE w/ PointGPT-B Ours 120.5 96.9($\uparrow$ 1.1) 95.6($\uparrow$ 0.4) 92.4($\uparrow$ 0.5) 94.7($\uparrow$ 0.3)
+Table 1: Classification accuracy on real-scanned (ScanObjectNN) and synthetic (ModelNet40) point clouds.
+In ScanObjectNN, we report the overall accuracy (%) on three variants.
+In ModelNet40, we report the overall accuracy (%) for both without and with voting.
+”#Params” represents the model’s parameter count.
+$\dagger$ indicates that the pre-trained model used simple rotational augmentation during fine-tuning on ScanObjectNN.
 4 Experiments
-First, we pre-train the PointHDMAE model using our block-to-scene pretraining strategy based on point cloud data from the scene domain.
+First, we pre-train the Point-MoDE model using our block-to-scene pretraining strategy based on point cloud data from the scene domain.
 After pre-training, we directly transfer the pre-trained model to various downstream tasks in different point cloud domains for fine-tuning.
-During fine-tuning, we selectively activate different sub-modules depending on the domain of the point cloud; for instance, we activate the object encoder for object point clouds, while utilizing all encoders for scene point clouds.
-This strategy enables our PointHDMAE model, pre-trained with the block-to-scene approach, to outperform existing domain-specific models in most cases without any additional domain adaptation training, demonstrating the generalization capability of our model.
-4.1 Pre-training
-Datasets.
-We combined all training data from the two most commonly used indoor scene datasets, SUNRGB-D (Song et al., 2015) and ScanNetV2 (Dai et al., 2017), to construct our pretraining dataset.
+During fine-tuning, we selectively activate different sub-modules depending on the domain of the point cloud; for instance, we activate the object expert encoder for object point clouds, while utilizing all encoders for scene point clouds.
+This strategy enables Point-MoDE model, pre-trained with the block-to-scene approach, to outperform existing domain-specific models in most cases without any additional domain adaptation training, demonstrating the superiority and transferability of our model.
+4.1 Dataset and Pre-training
+We initialize our object expert model using the Point-MAE or other models pre-trained on ShapeNet Chang et al.
+(2015), which can be directly obtained from these official repositories.
+Then, We combined all training data from the two most commonly used indoor scene datasets, SUNRGB-D Song et al.
+(2015) and ScanNetV2 Dai et al.
+(2017), to construct our block-to-scene pretraining dataset.
 Specifically, SUNRGB-D includes 5K single-view RGB-D training samples with oriented bounding box annotations for 37 object categories.
@@
 We extracted 50K points for each of the 6.2K samples, using only the xyz coordinates of each point to construct the pretraining dataset.
-Table 1: Classification accuracy on real-scanned (ScanObjectNN) and synthetic (ModelNet40) point clouds.
-In ScanObjectNN, we report the overall accuracy (%) on three variants.
-In ModelNet40, we report the overall accuracy (%) for both without and with voting.
-"#Params" represents the model’s parameter count.
-Method Reference #Params (M) ScanObjectNN ModelNet40
-OBJ-BG OBJ-ONLY PB-T50-RS w/o Vote w/ Vote
-Supervised Learning Only
-PointNe (Qi et al., 2017a) CVPR 2017 3.5 73.3 79.2 68 89.2 -
-PointNet++ (Qi et al., 2017b) NeurIPS 2017 1.7 82.3 84.3 77.9 90.7 -
-DGCNN (Wang et al., 2019) TOG 2019 1.8 82.8 86.2 78.1 92.9 -
-PointMLP (Ma et al., 2022) ICLR 2022 12.6 - - 85.2 94.1 94.5
-SFR (Zha et al., 2023a) ICASSP 2023 - - - 87.8 93.9 -
-P2P-HorNet (Wang et al., 2022) NeurIPS 2022 195.8 - - 89.3 94.0 -
-Single Modal Self-Supervised Learning
-Point-BERT (Yu et al., 2022) CVPR 2022 22.1 87.43 88.12 83.07 92.7 93.2
-MaskPoint (Liu et al., 2022) ECCV 2022 22.1 89.30 88.10 84.30 - 93.8
-Point-MAE (Pang et al., 2022) ECCV 2022 22.1 90.02 88.29 85.18 93.2 93.8
-Point-M2AE (Zhang et al., 2022) NeurIPS 2022 15.3 91.22 88.81 86.43 93.4 94.0
-PointGPT-S (Chen et al., 2024) NeurIPS 2023 29.2 91.63 90.02 86.88 - 94.0
-PointDif (Zheng et al., 2024) CVPR 2024 - 93.29 91.91 87.61 - -
-MaskFeat3D (Yan et al., 2024) ICLR 2024 15.3 93.20 91.50 88.40 - 94.0
-PointMamba (Liang et al., 2024) NeurIPS 2024 12.3 94.32 92.60 89.31 93.6 -
-PointHDMAE (Ours) Ours 22.8 95.18 93.12 90.01 93.7 94.2
-Multimodal Self-Supervised Learning
-Joint-MAE (Guo et al., 2023) IJCAI 2023 - 90.94 88.86 86.07 - 94.0
-ACT (Dong et al., 2023) ICLR 2023 22.1 93.29 91.91 88.21 93.2 93.7
-TAP+PointMLP (Wang et al., 2023) ICCV 2023 12.6 - - 88.50 94.0 -
-I2P-MAE (Zhang et al., 2023) CVPR 2023 15.3 94.15 91.57 90.11 93.7 94.1
-Recon (Qi et al., 2023) ICML 2023 44.3 95.18 93.29 90.63 94.1 94.5
-Pre-training.
-During the pretraining phase, we input all 50K×3 point clouds into the scene encoder of PointHDMAE to extract scene-level features.
+During the pretraining phase, we input all 50K×3 point clouds into the scene encoder of Point-MoDE to extract scene-level features.
 Simultaneously, we randomly select 32 point blocks from the scene point cloud, with each block containing 2K local points, and input these into the 32 object encoders with shared parameters.
-We use the AdamW (Kingma & Ba, 2014) optimizer with a base learning rate of 5e-4 and a weight decay of 0.1.
+We use the AdamW optimizer with a base learning rate of 5e-4 and a weight decay of 0.1.
 Simple rotation is applied as data augmentation to both the scene point cloud and each selected object point cloud.
-For the object point cloud masks, we set the mask ratio to 60% following previous work (Pang et al., 2022; Dong et al., 2023).
-We train the model from scratch for 200 epochs using 8 A100 GPUs.
-Furthermore, we can also leverage pre-trained object point cloud models on the ShapeNet55 (Chang et al., 2015) dataset to initialize our object-level models.
+For the object point cloud masks, we set the mask ratio to 60% following previous work Pang et al.
+(2022); Dong et al.
+(2023).
+We train the model for 200 epochs using 8 A100 GPUs.
 4.2 Fine-tuning on Downstream Tasks
 4.2.1 Object Point Cloud Classification
-We first evaluate the performance of our model on object point cloud classification tasks using the object encoder of PointHDMAE.
-We conduct point cloud classification on two of the most commonly used object point cloud datasets: ScanObjectNN (Uy et al., 2019) and ModelNet40 (Wu et al., 2015).
+We first evaluate the performance of our model on object point cloud classification tasks using the object expert encoder of Point-MoDE.
+We conduct point cloud classification on two of the most commonly used object point cloud datasets: ScanObjectNN Uy et al.
+(2019) and ModelNet40 Wu et al.
+(2015).
 ScanObjectNN contains 15K real scanned point clouds, each with various backgrounds, occlusions, and noise, which effectively assesses the model’s robustness.
 ModelNet40 includes 12K synthetic point clouds belonging to 40 different categories, with each point cloud being complete and clean, providing a better representation of 3D object shapes.
-Following previous work (Dong et al., 2023; Liang et al., 2024), we use 2K points as input for ScanObjectNN, apply simple rotation for data augmentation, and report classification accuracy without using voting.
-For ModelNet40, we use 1K points as input, apply scale and translate data augmentation, and report classification accuracy both without voting and with the standard voting mechanism.
-As presented in Table 1, firstly, compared to the recent state-of-the-art method PointMamba, our approach surpasses it by 0.86%, 0.52%, and 0.70% on the three variants of ScanObjectNN, respectively.
-This improvement is significant, given that the same downstream task settings were used.
-Secondly, our method surpasses the majority of multimodal pre-trained models, ranking just the same with the leading Recon (Qi et al., 2023).
-This is still highly competitive as Recon (Qi et al., 2023) benefits from the supplementary knowledge of image, and language modalities, while also requiring significantly more parameters than our method.
-Our PointHDMAE achieves leading performance indicating that using randomly selected point blocks from the scene, despite lacking explicit real-world significance, can still be effectively used for mask reconstruction pre-training.
-This effectiveness arises because mask reconstruction primarily focuses on learning general representations through the reconstruction of the original shapes, without requiring a specific understanding of the shapes’ concrete meanings.
-4.2.2 Scene Point Cloud Detection
-We further fine-tune the pre-trained PointHDMAE on scene-level object detection tasks.
-At this stage, we primarily rely on the scene encoder to process scene-level inputs.
-Simultaneously, we randomly select 32 point blocks from the scene point cloud and use the 32 shared object encoders to handle these local point blocks.
-During the decoding phase, we integrate the results from the scene encoder with random queries from the scene, helping the scene encoder to better understand scene details.
-We use ScanNetV2 (Dai et al., 2017), to evaluate our model’s scene understanding capabilities.
-Table 4.2.2 shows our detection results, our PointHDMAE model has shown significant improvements compared with other models.
-For example, compared to the previous state-of-the-art domain-specific pretraining model PointDif (Zheng et al., 2024), our method achieves a 6.2% improvement on $AP_{50}$.
-This substantial improvement is mainly attributed to our PointHDMAE using multiple object encoders to assist the scene encoder in analyzing the overall scene.
-This approach enables the scene model to focus on more local details, thereby enhancing scene understanding.
-Table 2: Object detection results on ScanNetV2. We adopt the average precision with 3D IoU thresholds of 0.25 ($AP_{25}$) and 0.5 ($AP_{50}$) for the evaluation metrics.
-Methods $AP_{25}$ $AP_{50}$ Supervised Learning Only VoteNet (Qi et al., 2019) 58.6 33.5 3DETR (Misra et al., 2021) 62.1 37.9 Single Modal Self-Supervised Learning PointContrast (Xie et al., 2020b) 58.5 38.0 STRL (Huang et al., 2021) 59.5 38.4 DepthContrast (Zhang et al., 2021) 61.3 - Point-BERT (Yu et al., 2022) 61.0 38.3 MaskPoint (Liu et al., 2022) 64.2 42.1 PointDif (Zheng et al., 2024) - 43.7 PointHDMAE (Ours) 66.8 49.9 Multimodal Self-Supervised Learning PiMAE (Chen et al., 2023) 62.6 39.4 ACT (Dong et al., 2023) 63.8 42.1 TAP (Wang et al., 2023) 62.6 39.4
-Table 3: Part segmentation results on the ShapeNetPart.
-The mean IoU across all categories, i.e., $\mathrm{mIoU}_{c}$ (%), and the mean IoU across all instances, i.e., $\mathrm{mIoU}_{I}$ (%) are reported.
-Methods $\mathrm{mIoU}_{c}$ $\mathrm{mIoU}_{I}$ Supervised Learning Only PointNet++ (Qi et al., 2017b) 81.9 85.1 PointMLP (Ma et al., 2022) 84.6 86.1 Single-Modal Self-Supervised Learning PointContrast (Xie et al., 2020b) - 85.1 CrossPoint (Afham et al., 2022) - 85.5 Point-BERT (Yu et al., 2022) 84.1 85.6 MaskPoint (Liu et al., 2022) 84.4 86.0 Point-MAE (Pang et al., 2022) 84.2 86.1 PointGPT-S (Chen et al., 2024) 84.1 86.2 PointMamba (Liang et al., 2024) 84.4 86.2 PointHDMAE (Ours) 85.0 86.3 Multimodal Self-Supervised Learning ACT (Dong et al., 2023) 84.7 86.1 Joint-MAE (Guo et al., 2023) 85.4 86.3 Recon (Qi et al., 2023) 84.8 86.4
+Following previous work Dong et al.
+(2023); Liang et al.
+(2024), we use 2K points as input for ScanObjectNN, apply simple rotation for data augmentation, and report classification accuracy without using voting.
+For ModelNet40, we use 1K points as input, apply scale and translate data augmentation, and report classification accuracy with the standard voting mechanism.
+As presented in Table 1, firstly, compared to previous single-domain baselines (Point-MAE, Point-BERT, and PointGPT), our Point-MoDE method achieves consistent and stable improvements, indicating that leveraging scene domain knowledge to assist object domain learning is an effective strategy.
+This approach not only preserves object representations but also enables the learning of more comprehensive knowledge.
+Secondly, our Point-MoDE shows a significant improvement on the real-world scanned point cloud dataset ScanObjectNN, while the improvement on the synthetic point cloud dataset ModelNet40 is smaller.
+This is because we use randomly sampled scene data during pretraining, which lacks explicit semantics.
+Such non-semantic data is more similar to real scanned point clouds, thereby aiding the model in learning more realistic point cloud representations.
 4.2.2 Object Point Cloud Part Segmentation
-We assess the performance of our PointHDMAE in part segmentation using the ShapeNetPart dataset (Chang et al., 2015), comprising 16,881 samples across 16 categories.
+We assess the performance of our Point-MoDE in part segmentation using the ShapeNetPart dataset Chang et al.
+(2015), comprising 16,881 samples across 16 categories.
 We utilize the same segmentation setting after the pre-trained encoder as in previous works Pang et al.
@@
 (2022) for fair comparison.
-The experimental results, displayed in Table 4.2.2, demonstrate that our model exhibits competitive performance in tasks such as part segmentation, which demands a more fine-grained understanding of point clouds.
-4.2.4 Object Point Cloud Completion
-Previous pretraining models have not explored the effects on low-level tasks.
-We first specifically design task heads for downstream low-level point cloud completion tasks.
-Then, we fine-tune our PointHDMAE model on point cloud completion on the classic point cloud completion dataset PCN (Yuan et al., 2018) and ShapeNet-55 (Chang et al., 2015).
-The PCN dataset is created from the ShapeNet (Chang et al., 2015) dataset, including eight categories with a total of 30974 CAD models.
-Compared to the PCN dataset, ShapeNet-55 includes 55 different categories of 3D models.
-Following previous practices (Yu et al., 2021; Li et al., 2023), we used 41,952 models for training and 10,518 models for testing.
-For each object, we randomly sampled 8,192 points from the surface to obtain the point cloud.
-We also divide the test samples into three difficulty degrees, simple, moderate, and hard in our experiments and we report the performance for each method in simple, moderate, and hard to show the ability of each network to deal with tasks at difficulty levels.
-We followed the data processing methods established in previous works (Yu et al., 2021; Li et al., 2023) and report the average $l_{1}$ Chamfer Distance (Fan et al., 2017) (CD-Avg) across all 8 object categories in the PCN dataset, and the average Chamfer Distance for the 55 categories under the simple (CD-S), moderate (CD-M), and hard (CD-H) settings in the ShapeNet-55 dataset, along with the overall average of these three metrics (CD-Avg).
-As shown in table 4, our model achieves state-of-the-art results across various settings in both datasets, demonstrating that our pre-trained model can better handle diverse scenarios, such as different viewpoints, object categories, incomplete patterns, and varying levels of incompleteness, even in lower-level completion tasks.
-Table 4: Quantitative comparison of point cloud completion task on PCN. Point resolutions for the output and ground-truth are 16384. For Chamfer Distance, lower is better.
-Methods Reference PCN ShapeNet-55
-CD-Avg CD-S CD-M CD-H CD-Avg
-FoldingNet (Yang et al., 2018) CVPR 2018 14.31 2.67 2.66 4.05 3.12
-PCN (Yuan et al., 2018) 3DV 2018 9.64 1.94 1.96 4.08 2.66
-GRNet (Xie et al., 2020a) ECCV 2020 8.83 1.35 1.71 2.85 1.97
-PoinTr (Yu et al., 2021) ICCV 2021 8.38 0.58 0.88 1.79 1.09
-LAKeNet (Tang et al., 2022) CVPR 2022 7.23 - - - 0.89
-SnowFlakeNet (Xiang et al., 2021) ICCV 2021 7.21 0.70 1.06 1.96 1.24
-ProxyFormer (Li et al., 2023) CVPR 2023 6.77 0.49 0.75 1.55 0.93
-SeedFormer (Zhou et al., 2022) ECCV 2022 6.74 0.50 0.77 1.49 0.92
-PointHDMAE Ours 6.54 0.51 0.70 1.24 0.81
-4.2.5 Scene Semantic Segmentation
-We have conducted experiments on scene-level semantic segmentation tasks to assess the performance of PointHDMAE in classifying each point in a scene into semantic categories.
-We validated our model using the indoor S3DIS (Armeni et al., 2016) dataset for semantic segmentation tasks.
+The experimental results displayed in Table 2.
+Our method also achieves consistent improvements compared to single-domain approaches, demonstrating the superiority of leveraging object knowledge to assist in scene tasks.
+This improvement appears less pronounced compared to detection tasks, which can be attributed to the inherently challenging nature of fine-grained tasks like segmentation.
+Therefore, even relatively small improvements are sufficient to highlight the effectiveness of our approach.
+Methods Reference $\mathrm{mIoU}_{c}$ $\mathrm{mIoU}_{I}$
+Supervised Learning Only
+PointNet++ Qi et al. (2017b) NeurIPS 2017 81.9 85.1
+PointMLP Ma et al. (2022) ICLR 2022 84.6 86.1
+Single-Modal Self-Supervised Learning
+PointContrast Xie et al. (2020) ECCV 2020 - 85.1
+IDPT Zha et al. (2023b) ICCV 2023 83.8 85.9
+Point-BERT Yu et al. (2022) CVPR 2022 84.1 85.6
+MaskPoint Liu et al. (2022) ECCV 2022 84.4 86.0
+Point-MAE Pang et al. (2022) ECCV 2022 84.2 86.1
+ACT Dong et al. (2023) ICLR 2023 84.7 86.1
+Point-FEMAE Zha et al. (2024a) AAAI 2024 84.9 86.3
+PointGPT-B Chen et al. (2024) NeurIPS 2023 84.5 86.5
+Point-MoDE w/ Point-BERT Ours 84.6($\uparrow$ 0.5) 86.2($\uparrow$ 0.6)
+Point-MoDE w/ Point-MAE Ours 84.5($\uparrow$ 0.3) 86.3($\uparrow$ 0.2)
+Point-MoDE w/ PointGPT-B Ours 84.9($\uparrow$ 0.4) 86.5(--)
+Table 2: Part segmentation results on the ShapeNetPart.
+The mean IoU across all categories, i.e., $\mathrm{mIoU}_{c}$ (%), and the mean IoU across all instances, i.e., $\mathrm{mIoU}_{I}$ (%) are reported.
+4.2.3 Scene Point Cloud Detection
+We further fine-tune the pre-trained Point-MoDE on scene-level object detection tasks.
+At this stage, we primarily rely on the scene expert encoder to process scene-level inputs.
+Simultaneously, we randomly select 32 point blocks from the scene point cloud and use the 32 shared object encoders to handle these local point blocks.
+During the decoding phase, we integrate the results from the scene expert encoder with random queries from the scene, helping the scene expert to better understand scene details.
+We use ScanNetV2 Dai et al.
+(2017), to evaluate our model’s scene understanding capabilities.
+Table 3 shows our detection results, our Point-MoDE model has shown significant improvements compared with other single domain models.
+In particular, there is an improvement of 6-9% in the $AP_{50}$ metric.
+This substantial improvement is mainly attributed to our Point-MoDE using multiple object encoders to assist the scene encoder in analyzing the overall scene.
+This approach enables the scene model to focus on more local details, thereby enhancing scene understanding.
+Methods Reference $AP_{25}$ $AP_{50}$
+Supervised Learning Only
+VoteNet Qi et al. (2019) ICCV 2019 58.6 33.5
+3DETR Misra et al. (2021) ICCV 2021 62.1 37.9
+Self-Supervised Learning
+PointContrast Xie et al. (2020) ECCV 2020 58.5 38.0
+DepthContrast Zhang et al. (2021) ICCV 2021 61.3 -
+Point-BERT Yu et al. (2022) CVPR 2022 61.0 38.3
+PiMAE Chen et al. (2023) CVPR 2023 62.6 39.4
+Point-MAE Pang et al. (2022) ECCV 2022 59.5 41.2
+ACT Dong et al. (2023) ICLR 2023 63.8 42.1
+MaskPoint Liu et al. (2022) ECCV 2022 64.2 42.1
+PointGPT-B Pang et al. (2022) NeurIPS 2023 63.1 42.8
+PointDif Zheng et al. (2024) CVPR 2024 - 43.7
+Point-MoDE w/ Point-BERT Ours 65.4($\uparrow$ 4.4) 47.1($\uparrow$ 8.8)
+Point-MoDE w/ Point-MAE Ours 66.2($\uparrow$ 6.7) 48.8($\uparrow$ 7.6)
+Point-MoDE w/ PointGPT-B Ours 66.7($\uparrow$ 3.6) 48.9($\uparrow$ 6.1)
+Table 3: Object detection results on ScanNetV2. We adopt the average precision with 3D IoU thresholds of 0.25 ($AP_{25}$) and 0.5 ($AP_{50}$) for the evaluation metrics.
+4.3 Scene Semantic Segmentation
+We have conducted experiments on scene-level semantic segmentation tasks to assess the performance of Point-MoDE in classifying each point in a scene into semantic categories.
+We validated our model using the indoor S3DIS Armeni et al.
+(2016) dataset for semantic segmentation tasks.
 Specifically, we tested the model on Area 5 while training on other areas and report the mean IoU (mIoU) and mean Accuracy (mAcc).
-To ensure a fair comparison, we used the same codebase based on the PointNext (Qian et al., 2022a) baseline and employed identical decoders and semantic segmentation heads.
-We acknowledge that there are several outstanding works in semantic segmentation, such as PointTransformerV3 (Wu et al., 2024), that achieve performance far exceeding that of PointNeXt.
-However, these advanced models often require more complex inputs, such as color and normal information.
-Our focus is on segmentation using only the most basic xyz coordinate information.
-Therefore, we chose PointNeXt as the baseline codebase for our pretraining model and ensure a fair comparison with other pretraining models.
-The experimental results are shown in the table 4.2.5.
+To ensure a fair comparison, we used the same codebase based on the PointNext Qian et al.
+(2022a) baseline and employed identical decoders and semantic segmentation heads.
+The experimental results are shown in the table 4.
 Compared to training the PointNeXt model from scratch, our method improves the mIoU score by 2.3%.
-It also shows significant improvements over other pretraining models, such as Point-MAE (Pang et al., 2022) and PointDif (Zheng et al., 2024).
+It also shows significant improvements over other pretraining models, such as Point-MAE Pang et al.
+(2022) and PointDif Zheng et al.
+(2024).
 This enhancement is largely due to our block-to-scene pretraining, which equips the model with strong scene understanding capabilities and further demonstrates the generalizability of our approach.
+Methods Reference mIoU mAcc
+Supervised Learning Only
+PointNeXt Qian et al. (2022a) NeurIPS 2022 68.5 75.1
+Pix4Point Qian et al. (2022b) 3DV 2024 69.6 75.2
+Self-Supervised Learning
+Point-BERT Yu et al. (2022) CVPR 2022 68.9 76.1
+MaskPoint Liu et al. (2022) ECCV 2022 68.6 74.2
+Point-MAE Pang et al. (2022) ECCV 2022 68.4 76.2
+PointDif Zheng et al. (2024) CVPR 2024 70.0 77.1
+PointGPT-B Pang et al. (2022) NeurIPS 2023 70.4 77.9
+Point-MoDE w/ Point-BERT Ours 69.3($\uparrow$ 0.4) 76.8($\uparrow$ 0.7)
+Point-MoDE w/ Point-MAE Ours 69.2($\uparrow$ 0.8) 76.6($\uparrow$ 0.4)
+Point-MoDE w/ PointGPT-B Ours 70.9($\uparrow$ 0.5) 78.1($\uparrow$ 0.2)
 Table 4: Semantic segmentation results on S3DIS Area 5.
-Methods mIoU mAcc PointCNN (Li et al., 2018) 57.3 63.9 Pix4Point (Qian et al., 2022b) 69.6 75.2 PointNeXt (Qian et al., 2022a) 68.5 75.1 Point-BERT (Yu et al., 2022) 68.9 76.1 MaskPoint (Liu et al., 2022) 68.6 74.2 Point-MAE (Pang et al., 2022) 68.4 76.2 PointDif (Zheng et al., 2024) 70.0 77.1 PointHDMAE (Ours) 70.8 77.6
-Table 6: Head tuning of object-level classification and scene-level detection.
-Object-Level Classification Methods OBJ-BG OBJ-ONLY Point-BERT (Yu et al., 2022) 88.81 88.3 Point-MAE (Pang et al., 2022) 89.50 88.98 PointHDMAE (Ours) 90.71 88.47 Scene-Level Detection Methods $AP_{25}$ $AP_{50}$ Point-BERT (Yu et al., 2022) 60.13 38.5 Point-MAE (Pang et al., 2022) 60.82 40.4 PointHDMAE (Ours) 62.44 41.8
-4.2.6 Head Tuning on Downstream Tasks
-We further demonstrate the generalization capability of our pre-trained model by performing only head tuning on downstream tasks.
-We keep the pre-trained feature extractor fixed and only train the task-specific heads (including the classification head and detection head).
-At the same time, all models use the same downstream task setting as described in 4.2.
-These experiments are designed to better evaluate the performance of the proposed PointHDMAE and to facilitate a more direct comparison with the baselines.
-The results of these experiments are summarized in the tables 4.2.5.
-As shown, our proposed Point-HDMAE still demonstrates superior performance compared to the baselines, highlighting its robust representational capabilities even without fine-tuning the entire model.
 4.4 Ablation Study
-4.3.1 The Impact of Stop Gradients.
-In our implementation, stopping gradients is crucial.
-This approach allows different encoders within the model to learn scene-level and object-level representations independently during block-to-scene pretraining.
-By ensuring that each encoder focuses solely on its specific task, we enhance the model’s generalization capability.
-Consequently, maintaining the accuracy and independence of each encoder’s learning objective during training is essential.
+4.4.1 The Impact of Different Pretraining Strategies
+Scene B. R. Object M. R. Coordinate S. T. Joint T. Classification Detection
+88.27 45.8
+88.32 41.2
+88.86 41.2
+88.32 43.8
+88.86 43.8
+87.21 47.4
+89.14 48.8
+Table 5: The Impact of Different Pretraining Strategies.
+We first analyze the impact of each key component in the block-to-scene pretraining strategy.
+Our block-to-scene pretraining strategy consists of four key parts: scene-level block position regression (Scene B.
+P.), object-level block mask reconstruction (Object M.
+R.), coordinate space transformation (Coordinate S.
+T.) of object blocks (from scene space to object space), and joint pertaining (Joint T.) of objects and scenes (connecting objects and scenes via the Projection module in Figure 3).
+A detailed analysis of these components is provided in Table 5.
+When using only Scene B.
+P.
+or Object M.
+R., the pretraining strategy enhances only the subsequent scene tasks or object tasks, failing to achieve simultaneous improvements.
+Even when both strategies are employed without Joint T., while simultaneous enhancement is possible, the inability of the two pipelines to interact prevents the complementary knowledge from being effectively learned, thereby limiting the representation capability.
+Only by jointly employing both regression and reconstruction can the most comprehensive representation be achieved.
+Secondly, Coordinate S.
+T.
+is crucial for learning object representations.
+This is because the coordinates of blocks directly cropped from the scene belong to the scene space and are often not normalized within the range of [-1, 1].
+In contrast, most object-level point cloud analysis assumes coordinates within the range of [-1, 1].
+Thus, transforming the coordinates is essential.
+As shown in the table, omitting this transformation results in performance degradation on object tasks.
+This is because the scene-space coordinates disrupt the learning of object representations, highlighting the importance of this normalization step for effective object-level analysis.
+4.4.2 The Impact of Stop Gradients
+Methods Object Classification Scene Detection
+ModelNet40 ScanObjectNN AP25 AP50
+w/o S. G. 93.4 87.45 62.7 41.5
+w/ S. G. 94.0 89.14 66.2 48.8
+Table 6: The impact of stop gradients.
+In our implementation, gradient stopping plays a crucial role.
+This mechanism allows different encoders within the model to jointly learn domain-specific representations during block-to-scene pretraining without interfering with each other.
+Thus, ensuring the accuracy and independence of each encoder’s learning objectives during training is of paramount importance.
 Since the position regression objective and the masked reconstruction objective are two distinct tasks in our pretraining process, failing to decouple the learning processes of the different encoders could lead to catastrophic forgetting.
@@
 By applying gradient stopping, we effectively prevent this interference, ensuring that each encoder remains focused on its specific task and thereby avoiding catastrophic forgetting.
-To further validate this issue, we conducted experiments where we removed the Stop Gradients operation from our pipeline and retrained the PointHDMAE model.
-We then assessed its performance on both scene-level point cloud detection tasks and object-level classification tasks.
-This comparison allowed us to observe the impact of gradient stopping on the model’s ability to effectively learn and generalize across different tasks.
-As shown in table 4.3.1, after removing the Stop Gradients operation, there was a 8.6% decrease in AP50 for the scene-level detection tasks, clearly indicating a deterioration in the representation capabilities of the scene encoder.
-Similarly, there was a noticeable decline in performance across classification tasks.
-These results collectively demonstrate that decoupling the representation learning of different encoders is essential to avoid catastrophic forgetting.
-Table 7: The impact of stop gradients.
-Object-Level Classification OBJ-BG OBJ-ONLY w/o Stop Gradients 93.29 92.43 w/ Stop Gradients 95.18 93.12 Scene-Level Detection $AP_{25}$ $AP_{50}$ w/o Stop Gradients 63.1 41.3 w/ Stop Gradients 66.8 49.9
-Table 8: The impact of the number of point blocks.
-Number of blocks FLOPs(G) PB-T50-RS 1 22 88.78 8 59 89.10 16 102 89.52 32 187 90.01 64 359 90.12
-4.3.2 The Impact of the Number of Point Blocks.
-The number of randomly selected point blocks during the pretraining phase has a significant impact on the representation learning of the object encoder.
-Each point block serves as a sample to train the object encoder.
-The more point blocks selected from a scene, the richer the knowledge the object encoder can learn.
-However, this also leads to a significant increase in computational complexity.
-Therefore, we need to select an appropriate number to achieve a balance between efficiency and performance.
-We selected 1, 8, 16, 32, and 64 blocks respectively and pre-trained these models from scratch.
-We reported the computational floating-point operations (FLOPs) required for pre-training and the performance of the resulting object models.
-The trained models were fine-tuned on the PB-RS-T50 variant of ScanObjectNN, using its split to evaluate model performance.
-As illustrated in the table 4.3.1, our model’s computational complexity significantly increases with the number of blocks, while the performance of the model gradually improves and eventually levels off.
-To achieve a balance between performance and efficiency, we chose 32 blocks for our experiments.
-4.4 Conclusion
-In this paper, we first propose a point cloud hybrid-domain masked autoencoders model to address the generalization limitations of existing domain-specific models.
-Our hybrid model selectively activates the object encoder to handle object domain point clouds specifically and leverages these object encoders to assist the scene encoder in processing scene domain point clouds.
-Then, we propose a block-to-scene pre-training strategy to train our PointHDMAE model.
-This strategy involves joint training through random object mask reconstruction and position regression within the scene, enabling our domain-specific encoder models to learn general representations relevant to their respective domains.
-Finally, we demonstrated the generalization and superiority of our model through extensive experiments across various datasets and tasks from different domains.
+Table 6 reports the impact of applying gradient stopping.
+When the gradient stopping mechanism is removed, the AP50 on scene-level detection tasks drops by 7.3%, clearly indicating that the learning process of the scene encoder is disrupted.
+Similarly, significant performance degradation is observed in classification tasks.
+These results collectively demonstrate that decoupling the representation learning of different encoders is essential for preventing catastrophic forgetting.
+4.4.3 The Impact of Object Expert Initialization
+Methods Object Classification Scene Detection
+ModelNet40 ScanObjectNN AP25 AP50
+Scratch 93.2 88.38 65.7 48.5
+Pre-trained Point-MAE 94.0 89.14 66.2 48.8
+Table 7: The impact of object expert initialization.
+We further analyze the impact of initializing the object expert model.
+In our implementation, we initialize the model using an object model pre-trained on ShapeNet, after which no object data is used.
+This initialization ensures that we can learn representations from the object domain prior.
+We further examine the effect of not using object priors (Scratch), where no object domain data is used during training.
+The results, as shown in Table 7, indicate that the absence of object priors has a minimal impact on scene representation learning, resulting in only slight performance degradation.
+However, it significantly affects the learning of object point cloud representations.
+For the real point cloud ScanObjectNN, even when trained from scratch, the model eventually surpasses the one trained only on ShapeNet55 (88.27).
+However, for the synthetic point cloud ModelNet40, there is a larger gap compared to the model trained only on ShapeNet55 (93.8).
+This is because the point blocks cut from the scene lack specific semantics, making their distribution more similar to real point clouds but significantly different from the distribution of synthetic point clouds, leading to poorer performance on synthetic point clouds.
+5 Conclusion
+In this paper, we propose a block-to-scene pretraining strategy to train a mixture-of-domain-experts model that seamlessly integrates scene domain and object domain experts.
+This strategy enables the simultaneous acquisition of object-level and scene-level representations, enabling the learning of comprehensive 3D representations.
+Extensive experiments conducted across various datasets and tasks validate the superiority and transferability of our proposed approach.
+Our findings highlight the potential of leveraging cross-domain knowledge to achieve robust and versatile 3D representation learning, paving the way for future study.
+Acknowledgements
+This work is supported in part by the National Natural Science Foundation of China, under Grant (62302309,62171248), Shenzhen Science and Technology Program (JCYJ20220818101014030,JCYJ20220818101012025), and the PCNL KEY project (PCL2023AS6-1).